Ten years ago your friend invested in a plot of land with the idea of eventually building a summer cottage on it. His plan was to place a "nest egg" of $12,000 in a long term Certificate of Deposit which paid 8% interest over a period of 20 years, compounded monthly. At the end of the 20 years, he would use the money to build the cottage.
The problem was the plot had 8 prairie dogs on it 10 years ago when he acquired the ground.
This year he made a count of the prairie dogs and found there were 102 on the same plot! A biologist friend assured him that the prairie dogs were growing exponentially according to the law of natural growth, and something needed to be done.

1. Write the formula for natural exponential growth for this particular prairie dog population.
Hint: Find the "k" value for this population, and put it in your formula.

2. Once you have this formula, predict:
a. approximately how many years from now will the population reach 365 prairie dogs.
b. approximately how large the population will be 10 years from now.

3. Write the formula for compound interest and calculate how much money your friend will have on deposit (in the account described above) after 20 years (10 years from now).

Answer by stanbon(75887)

The problem was the plot had 8 prairie dogs on it 10 years ago when he acquired the ground.
This year he made a count of the prairie dogs and found there were 102 on the same plot! A biologist friend assured him that the prairie dogs were growing exponentially according to the law of natural growth, and something needed to be done.
1. Write the formula for natural exponential growth for this particular prairie dog population.
You have two points: (-10,8) and (0,102)
x is years from now ; y = # of prairie dogs.
-------------------
Exponential Form:
y = ab^x
Substitute the points into the form:
1st: 8 = ab^-10
2nd: 102 = ab^0
--------------
From the 2nd equation you get a = 102
Substitute that into the 1st equation and solve for "b".
8 = 102b^-10
4/51 = b^-10
b^10 = 5/4
b = 1.022565
EQUATION:
y = 102(1.022565..)^x
-----------------
==========================
2. Once you have this formula, predict:
a. approximately how many years from now will the population reach 365 prairie dogs.
365 = 102(1.022565...)^x
(365/102) = 1.022565...^x
x = log(365/102)/log(1.022565...)
x = 2.5673 years
=========================
b. approximately how large the population will be 10 years from now.
Let x = 10 in the equation and solve for y.
---------------------------
Ten years ago your friend invested in a plot of land with the idea of eventually building a summer cottage on it. His plan was to place a "nest egg" of $12,000 in a long term Certificate of Deposit which paid 8% interest over a period of 20 years, compounded monthly. At the end of the 20 years, he would use the money to build the cottage.

3. Write the formula for compound interest and calculate how much money your friend will have on deposit (in the account described above) after 20 years (10 years from now).
A(t) = A(0)(1+r/n)^(nt)
A(20) = 12000(1+0.08/12)*(12*20)
A(20) = $59121,63
------------------
=====================
Cheers,
Stan H.